-4(t-9)-(t+4)=7 what is t equal to???
Found 2 solutions by stanbon, MathLover1:
Answer by stanbon(75887)

-4(t-9)-(t+4)=7 what is t equal to???
-------
-4t+36-t-4 = 7
-5t+32 = 7
-5t = -25
t = 5
=============
Cheers,
Stan H.

Answer by MathLover1(20855)